Question 15:

A firm is considering a project that has the following cash flow and cost of capital (r) data.

r =

11.00%

Year

0

1

2

3

Cash flows

−$800

$350

$350

$350

What is the project's MIRR?
Answer #1

Solution :

MIRR = ( Future value of cash inflow / present value of cash outlay )^ (1/year) - 1

Future value of cash flow = 350 *(1+11%)^2 + 350 *(1+11%)^1 + 350 * ( 1+11%) ^0

Future value of cash flow = 350 *1.11^2 + 350 * 1.11 + 350

Future value of cash flow = 350 *1.21 + 350 *1.1 + 350

Future value of cash flow = 423.5 + 385 + 350 = 1158.5

MIRR = ( 1158.5/800) ^(1/3) - 1 = 1.448125^(1/3) - 1 = 1.13136 -1 = 13.14%

MIRR = 13.14 %